# Transversal Hamilton cycles in digraph collections

### v1

@@ 5 Extremal case
Applying Claim 5.5 to the digraph collection $\{D_{i}^{\pm}[C^{2}\setminus V(P\cup Q_{0}),C^{4}\setminus V(P\cup Q_{0})]:i\in\mathcal{C}_{3}\setminus({\rm col}(P\cup Q_{0})\cup\{i_{1},i_{2},c_{1},c_{2}\})\}$, there exist two disjoint rainbow directed paths $P^{3}=x_{3}\ldots y_{3}$ and $P^{4}=x_{4}\ldots y_{4}$ with $x_{3}\in N_{{c_{1}}}^{+}(y_{2},C^{4}\setminus V_{\rm bad})$, $y_{3}\in N_{{i_{1}}}^{-}(u_{0},C^{4}\setminus V_{\rm bad})$, $x_{4}\in N_{{i_{2}}}^{+}(v_{0},C^{2}\setminus V_{\rm bad})$ and $y_{4}\in N_{{c_{2}}}^{-}(x_{1},C^{4}\setminus V_{\rm bad})$, respectively. Hence $\mathcal{D}$ contains a transversal directed Hamilton cycle, a contradiction.
@@
Now, assume that there exists a color $c\in\mathcal{C}_{\rm bad}\setminus{\rm col}(M\cup\mathbf{P}^{0})$. Together with the condition $\delta^{0}(\mathcal{D})\geq\frac{n}{2}$, there exists an edge $uv\in E(D_{c}[C^{2}\setminus V(Q_{0}\cup Q_{1}\cup Q_{2}),C^{3}\cup C^{4}])$ if $|C^{1}|+|C^{2}|\leq|C^{3}|+|C^{4}|$, and $vu\in E(D_{c}[C^{4}\setminus V(Q_{0}\cup Q_{1}\cup Q_{2}),C^{1}\cup C^{2}])$ otherwise. By a similar discussion as above and considering whether the vertex $v$ belongs to $V_{\rm bad}$ or not, we obtain that $\mathcal{D}$ contains a transversal directed Hamilton cycle, a contradiction.
@@
$\bullet$ $\sigma=1$, i.e., $|C^{4}\setminus V(P^{0})|=|C^{2}\setminus V(P^{0})|-1$. We first consider that there exists a color $c\in\mathcal{C}_{\rm bad}\setminus{\rm col}(M\cup\mathbf{P}^{0})$. Choose a directed edge, say ${xy}$, in $E(D_{c}[C^{3}\setminus V(Q_{0}\cup Q_{1}\cup Q_{3}),C^{1}\setminus V(Q_{0}\cup Q_{1}\cup Q_{3})])$. By applying Claim 5.5 to the digraph collection $\{D_{i}[C^{1}\setminus(V(Q_{0}\cup Q_{1}\cup Q_{2})\cup\{x\})]\cup D_{i}[C^{3}\setminus(V(Q_{0}\cup Q_{1}\cup Q_{2})\cup\{y\})]:i\in\mathcal{C}_{3}\setminus({\rm col}(Q_{0}\cup Q_{1}\cup Q_{2})\cup\{i_{1},i_{2},i_{3},i_{4},c\})\}$, we obtain two disjoint rainbow directed paths $P^{1}=x_{1}\ldots y_{1}$ and $P^{2}=x_{2}\ldots y_{2}$, with $y_{1}\in N_{i_{3}}^{-}(u_{1},C^{1})$, $x_{2}\in N_{i_{4}}^{+}(v_{2},C^{3})$, $x_{1}\in C_{c_{1}}^{1}\setminus V_{\rm bad}$ and $y_{2}\in C_{c_{2}}^{3}\setminus V_{\rm bad}$ for two unused colors $c_{1},c_{2}\in\mathcal{C}_{3}$. By connecting $Q_{1},{xy},Q_{3}$ in turn using Claim 5.15 (iv), we obtain a rainbow directed path $P=x_{1}\ldots y_{2}$ such that $V_{\rm bad}\setminus X^{\prime}\subseteq V(P)$, $\mathcal{C}_{\rm bad}\subseteq{\rm col}(P)$ and $|C^{4}\setminus V(P)|=|C^{2}\setminus V(P)|+1$. Similar to the previous discussion, $\mathcal{D}$ contains a transversal directed Hamilton cycle, a contradiction.
@@
In what follows, it suffices to consider the case that $|E(Q_{1})|+\cdots+|E(Q_{t})|+k<|C^{2}|-|C^{4}|-\sigma=r-s_{1}-2s_{2}-\sigma$. Move all vertices inside $Q_{1},\ldots,Q_{t}$, all vertices in $(V(Q_{1}^{\prime})\cup\ldots\cup V(Q_{k}^{\prime}))\cap C^{2}$ along with their in-neighbors in $C^{3}$ to $C^{4}$. Denote the resulting partition by $\tilde{C}^{1}\cup\tilde{C}^{2}\cup\tilde{C}^{3}\cup\tilde{C}^{4}$. Let $\tilde{\mathcal{C}}:=\mathcal{C}\setminus({\rm col}(\bigcup_{i\in[t]}Q_{i})\cup{\rm col}(\bigcup_{i\in[k]}Q_{i}^{\prime}[C^{3},C^{2}]))$. Therefore,
@@
which implies that $|\tilde{C}^{2}|\leq|\tilde{C}^{4}|$. We complete our proof by the subsequent lemma (i.e., Lemma 5.16), which leads to a contradiction by showing that $\mathcal{D}$ contains a transversal directed Hamilton cycle. ∎
@@
$=(|C^{4}|-\delta^{\frac{1}{2}}|Y_{2}|+|C^{3}|)|C^{2}||\mathcal{C}^{\prime}|.$ (5)
Thus, $|Y_{2}|\leq 3\sqrt{\delta}n$. Similarly, $|Y_{1}|\leq 3\sqrt{\delta}n$. Notice that Claim 5.5 and Claim 5.15 hold by setting $V_{\rm bad}:=Y$ and $\mathcal{C}_{3}:=\mathcal{C}^{\prime}$.
@@
Since $\delta^{0}(\mathcal{D})\geq\frac{n}{2}$, we have $|E(D_{i}[C^{2},C^{3}\cup C^{4}])|,|E(D_{i}[C^{1}\cup C^{4},C^{2}])|\geq|C^{2}|\left\lceil\frac{n}{2}\right\rceil$ for each $i\in\mathcal{C}^{\prime}$. This implies that for each $i\in\mathcal{C}^{\prime}$, there are at most $|C^{2}|(\left\lceil\frac{n}{2}\right\rceil+\gamma_{1}n)-|C^{2}|\left\lceil\frac{n}{2}\right\rceil=|C^{2}|\gamma_{1}n$ non-edges in $D_{i}[C^{2},C^{3}\cup C^{4}]$ and at most $|C^{2}|\gamma_{2}n$ non-edges in $D_{i}[C^{1}\cup C^{4},C^{2}]$. Therefore, Hence
@@
We proceed our proof by considering the values of $|Y_{1}|$ and $|Y_{2}|$.
@@
In view of Lemma 5.13, there exists a subset $Y_{1}^{\prime}\subseteq Y_{1}$ with size $\gamma_{1}n$ such that vertices in $Y_{1}\setminus Y_{1}^{\prime}$ can be covered by disjoint rainbow directed $P_{3}$ inside $\{D_{i}[Y_{1},C^{2}]:i\in\mathcal{C}^{\prime}\}$ with centers in $Y_{1}\setminus Y_{1}^{\prime}$; and a subset $Y_{2}^{\prime}\subseteq Y_{2}$ with size $\gamma_{2}n$ such that vertices in $Y_{2}\setminus Y_{2}^{\prime}$ can be covered by disjoint rainbow $P_{3}$ inside $\{D_{i}[C^{2},Y_{2}]:i\in\mathcal{C}^{\prime}\}$ with centers in $Y_{2}\setminus Y_{2}^{\prime}$. Therefore, by using colors in $\mathcal{C}^{\prime}$,
  • vertices in $(Y_{1}\cup Y_{2})\setminus(Y_{1}^{\prime}\cup Y_{2}^{\prime})$ can be covered by a set of disjoint rainbow directed $P_{3}$ with centers in $(Y_{1}\cup Y_{2})\setminus(Y_{1}^{\prime}\cup Y_{2}^{\prime})$ and endpoints in $C^{2}$,
  • vertices in $Y_{2}^{\prime}\setminus Y_{1}^{\prime}$ can be covered by a set of disjoint rainbow directed $P_{3}$ or $P_{4}$ with centers in $Y_{2}^{\prime}\setminus Y_{1}^{\prime}$, starting at $C^{4}$ and ending at $C^{2}$,
  • vertices in $Y_{1}^{\prime}\setminus Y_{2}^{\prime}$ can be covered by a set of disjoint rainbow directed $P_{3}$ or $P_{4}$ with centers in $Y_{1}^{\prime}\setminus Y_{2}^{\prime}$, starting at $C^{2}$ and ending at $C^{4}$,
  • vertices in $Y_{1}^{\prime}\cap Y_{2}^{\prime}$ can be covered by a set of disjoint rainbow directed $P_{3}$, $P_{4}$ or $P_{5}$ with centers in $Y_{1}^{\prime}\cap Y_{2}^{\prime}$ and endpoints in $C^{4}$.
@@
It is routine to verify that there exists a rainbow matching inside $D_{i}^{\pm}[C^{2}\setminus V(\mathbf{P}),C^{4}\setminus V(\mathbf{P})]\cup D_{i}[C^{1}]\cup D_{i}[C^{3}]\cup(\bigcup_{k\in[4]}D_{i}[C^{k},C^{k+1}]):i\in\mathcal{C}^{\prime\prime}\}$, say $M$. For all but at most one unused colors in $\mathcal{C}_{\rm bad}$, we can find a rainbow directed $P_{3}$ with center in $Y$ and endpoints in $Z$, where $\{Y,Z\}=\{C^{1},C^{3}\}$. If some color $c\in\mathcal{C}_{\rm bad}$ remains unused, then $D_{c}[C^{1},C^{3}]$ is nearly a complete bipartite graph and $D_{c}[C^{2}],D_{c}[C^{4}]$ are almost complete graphs. Then, by a similar argument to that in Theorem 5.14, $\mathcal{D}$ contains a transversal directed Hamilton cycle.
@@
We only consider the case that $|Y_{1}|\geq\gamma_{1}n$ and $|Y_{2}|<\gamma_{2}n$; other cases can be discussed similarly. Notice that vertices in $Y_{1}$ can be covered by a set of disjoint rainbow directed paths with centers in $Y_{1}$ and ending at $C^{4}\setminus(Y_{1}\cup Y_{2})$. In particular, those rainbow directed paths can be chosen such that none of them contains edges inside $\mathcal{D}[C^{4},C^{3}]$ or $\mathcal{D}[C^{1},C^{4}]$. Furthermore, vertices in $Y_{2}\setminus Y_{1}$ can be covered by a set of disjoint rainbow directed $P_{3}$ with centers in $Y_{2}\setminus Y_{1}$ and starting at $C^{4}\setminus Y$, using colors in $\mathcal{C}^{\prime}$.
In the digraph collection $\mathcal{D}[C^{4}\setminus Y]$, extend those directed rainbow paths or choose other disjoint directed rainbow directed paths into a set of disjoint maximal rainbow directed paths. Let $\mathbf{P}:=\{Q_{1},Q_{2},\ldots,Q_{t}\}$ be a set consisting of all disjoint rainbow paths in the above, each of which has at least one endpoint in $C^{4}\setminus Y$ and length $s_{i}\ (1\leq i\leq t)$ inside $C^{4}$. Let $\{e_{1},e_{2},\ldots,e_{r}\}$ be a set consisting of all disjoint rainbow edges in $\mathcal{D}[C^{4}\setminus Y,C^{3}]$, and let $\{\tilde{Q}_{1},\ldots,\tilde{Q}_{k}\}$ be a set of disjoint maximal directed rainbow $P_{3}$ in $\mathcal{D}^{\pm}[C^{4}\setminus Y,C^{1}]$ with centers in $C^{4}$ and endpoints in $C^{1}$, each of which is chosen while avoiding previously used vertices and colors.
@@
In this subcase, $\sum_{i=1}^{t}s_{i}+k+r<2\gamma n$. Assume that $V(Q_{i}[C^{4}]):=\{v_{s_{1}+\cdots+s_{i-1}+i},\ldots,v_{s_{1}+\cdots+s_{i}+i}\}$ for each $i\in[t]$. Let $w$ be an unused vertex in $C^{4}$ and $c_{1},c_{2}$ be two unused colors in $\mathcal{C}$. By the maximality of $\mathbf{P}$, in $D_{c_{1}}$ (resp. $D_{c_{2}}$), $w$ cannot be adjacent to the start vertices of $Q_{i}[C^{4}]$ (resp. adjacent from the end vertices of $Q_{i}[C^{4}]$) for all $i\in[t]$.

### v2

@@ 1. $(\dagger)$
Applying Lemma 6.2 to the digraph collection $\{D_{i}^{\pm}[C^{2}\setminus V(P\cup Q_{0}),C^{4}\setminus V(P\cup Q_{0})]:i\in\mathcal{C}_{3}\setminus({\rm col}(P\cup Q_{0})\cup\{i_{1},i_{2},c_{1},c_{2}\})\}$, we obtain two disjoint rainbow directed paths $P^{3}=x_{3}\ldots y_{3}$ and $P^{4}=x_{4}\ldots y_{4}$, where $x_{3}\in N_{{c_{1}}}^{+}(y_{2},C^{4}\setminus V_{\rm bad})$, $y_{3}\in N_{{i_{1}}}^{-}(u_{0},C^{4}\setminus V_{\rm bad})$, $x_{4}\in N_{{i_{2}}}^{+}(v_{0},C^{2}\setminus V_{\rm bad})$ and $y_{4}\in N_{{c_{2}}}^{-}(x_{1},C^{4}\setminus V_{\rm bad})$, respectively. Hence $\mathcal{D}$ contains a transversal directed Hamilton cycle, a contradiction.
@@
Now, assume that there exists a color $c\in\mathcal{C}_{\rm bad}\setminus{\rm col}(M\cup\mathbf{P}^{0})$. Since $\delta^{0}(\mathcal{D})\geq\frac{n}{2}$, there exists an edge $uv\in E(D_{c}[C^{2}\setminus V(Q_{0}\cup Q_{1}\cup Q_{2}),C^{3}\cup C^{4}])$ if $|C^{1}|+|C^{2}|\leq|C^{3}|+|C^{4}|$, and $vu\in E(D_{c}[C^{4}\setminus V(Q_{0}\cup Q_{1}\cup Q_{2}),C^{1}\cup C^{2}])$ otherwise. By an argument analogous to that above, and considering whether $v\in V_{\rm bad}$, we obtain that $\mathcal{D}$ contains a transversal directed Hamilton cycle, a contradiction.
@@
$\bullet$ Suppose that $\sigma=1$, i.e., $|C^{4}\setminus V(P^{0})|=|C^{2}\setminus V(P^{0})|-1$. We first consider that there exists a color $c\in\mathcal{C}_{\rm bad}\setminus{\rm col}(M\cup\mathbf{P}^{0})$. Choose a directed edge, say xy{xy}, in $E(D_{c}[C^{3}\setminus V(Q_{0}\cup Q_{1}\cup Q_{3}),C^{1}\setminus V(Q_{0}\cup Q_{1}\cup Q_{3})])$. By applying Lemma 6.2 to the digraph collection $\{D_{i}[C^{1}\setminus(V(Q_{0}\cup Q_{1}\cup Q_{2})\cup\{x\})]\cup D_{i}[C^{3}\setminus(V(Q_{0}\cup Q_{1}\cup Q_{2})\cup\{y\})]:i\in\mathcal{C}_{3}\setminus({\rm col}(Q_{0}\cup Q_{1}\cup Q_{2})\cup\{i_{1},i_{2},i_{3},i_{4},c\})\}$, we obtain two disjoint rainbow directed paths $P^{1}=x_{1}\ldots y_{1}$ and $P^{2}=x_{2}\ldots y_{2}$, where $y_{1}\in N_{i_{3}}^{-}(u_{1},C^{1})$, $x_{2}\in N_{i_{4}}^{+}(v_{2},C^{3})$, $x_{1}\in C_{c_{1}}^{1}\setminus V_{\rm bad}$ and $y_{2}\in C_{c_{2}}^{3}\setminus V_{\rm bad}$ for two unused colors $c_{1},c_{2}\in\mathcal{C}_{3}$. By sequentially connecting $Q_{1},{xy},Q_{3}$ via Claim 7.1 (iv), we obtain a rainbow directed path $P=x_{1}\ldots y_{2}$ such that $V_{\rm bad}\setminus X^{\prime}\subseteq V(P)$, $\mathcal{C}_{\rm bad}\subseteq{\rm col}(P)$ and $|C^{4}\setminus V(P)|=|C^{2}\setminus V(P)|+1$. Similar to the previous discussion, $\mathcal{D}$ contains a transversal directed Hamilton cycle, a contradiction.
@@
On the other hand, we consider that $|E(Q_{1})|+\cdots+|E(Q_{t})|+k<|C^{2}|-|C^{4}|-\sigma=r-s_{1}-2s_{2}-\sigma$. In this case, we move all vertices inside $Q_{1},\ldots,Q_{t}$, all vertices in $(V(Q_{1}^{\prime})\cup\ldots\cup V(Q_{k}^{\prime}))\cap C^{2}$ along with their in-neighbors in $C^{3}$ to $C^{4}$. Denote the resulting partition by $\tilde{C}^{1}\cup\tilde{C}^{2}\cup\tilde{C}^{3}\cup\tilde{C}^{4}$. Let $\tilde{\mathcal{C}}:=\mathcal{C}\setminus({\rm col}(\bigcup_{i\in[t]}Q_{i})\cup{\rm col}(\bigcup_{i\in[k]}Q_{i}^{\prime}[C^{3},C^{2}]))$. Therefore,
@@
which implies that $|\tilde{C}^{2}|\leq|\tilde{C}^{4}|$. We complete our proof by using the subsequent lemma (i.e., Lemma 7.2), which leads to a contradiction by showing that $\mathcal{D}$ contains a transversal directed Hamilton cycle. ∎
@@
$=(|C^{4}|-\delta^{\frac{1}{2}}|Y_{2}|+|C^{3}|)|C^{2}||\mathcal{C}^{\prime}|.$
Thus, $|Y_{2}|\leq 3\sqrt{\delta}n$. Similarly, $|Y_{1}|\leq 3\sqrt{\delta}n$. Notice that Lemma 6.2 and Claim 7.1 hold by setting $V_{\rm bad}:=Y$ and $\mathcal{C}_{3}:=\mathcal{C}^{\prime}$.
@@
Since $\delta^{0}(\mathcal{D})\geq\frac{n}{2}$, we have $|E(D_{i}[C^{2},C^{3}\cup C^{4}])|,|E(D_{i}[C^{1}\cup C^{4},C^{2}])|\geq|C^{2}|\left\lceil\frac{n}{2}\right\rceil$ for each $i\in\mathcal{C}^{\prime}$. This implies that for each $i\in\mathcal{C}^{\prime}$, there are at most $|C^{2}|(\left\lceil\frac{n}{2}\right\rceil+\gamma_{1}n)-|C^{2}|\left\lceil\frac{n}{2}\right\rceil=|C^{2}|\gamma_{1}n$ non-edges in $D_{i}[C^{2},C^{3}\cup C^{4}]$ and at most $|C^{2}|\gamma_{2}n$ non-edges in $D_{i}[C^{1}\cup C^{4},C^{2}]$. Therefore,
@@
We proceed with the proof by considering the values of $|Y_{1}|$ and $|Y_{2}|$.
@@
In view of Lemma 6.4, there exists a subset $Y_{1}^{\prime}\subseteq Y_{1}$ with size $\gamma_{1}n$ such that vertices in $Y_{1}\setminus Y_{1}^{\prime}$ can be covered by disjoint rainbow directed $P_{3}$ copies inside $\{D_{i}[Y_{1},C^{2}]:i\in\mathcal{C}^{\prime}\}$ with centers in $Y_{1}\setminus Y_{1}^{\prime}$; and a subset $Y_{2}^{\prime}\subseteq Y_{2}$ with size $\gamma_{2}n$ such that vertices in $Y_{2}\setminus Y_{2}^{\prime}$ can be covered by disjoint rainbow directed $P_{3}$ copies inside $\{D_{i}[C^{2},Y_{2}]:i\in\mathcal{C}^{\prime}\}$ with centers in $Y_{2}\setminus Y_{2}^{\prime}$. Therefore, by using colors in $\mathcal{C}^{\prime}$,
  • vertices in $(Y_{1}\cup Y_{2})\setminus(Y_{1}^{\prime}\cup Y_{2}^{\prime})$ can be covered by a set of disjoint rainbow directed $P_{3}$ copies with centers in $(Y_{1}\cup Y_{2})\setminus(Y_{1}^{\prime}\cup Y_{2}^{\prime})$ and endpoints in $C^{2}$,
  • vertices in $Y_{2}^{\prime}\setminus Y_{1}^{\prime}$ can be covered by a set of disjoint rainbow directed $P_{3}$ or $P_{4}$ copies with centers in $Y_{2}^{\prime}\setminus Y_{1}^{\prime}$, starting at $C^{4}$ and ending at $C^{2}$,
  • vertices in $Y_{1}^{\prime}\setminus Y_{2}^{\prime}$ can be covered by a set of disjoint rainbow directed $P_{3}$ or $P_{4}$ copies with centers in $Y_{1}^{\prime}\setminus Y_{2}^{\prime}$, starting at $C^{2}$ and ending at $C^{4}$,
  • vertices in $Y_{1}^{\prime}\cap Y_{2}^{\prime}$ can be covered by a set of disjoint rainbow directed $P_{3}$, $P_{4}$ or $P_{5}$ copies with centers in $Y_{1}^{\prime}\cap Y_{2}^{\prime}$ and endpoints in $C^{4}$.
@@
It is routine to verify that there exists a rainbow matching inside $D_{i}^{\pm}[C^{2}\setminus V(\mathbf{P}),C^{4}\setminus V(\mathbf{P})]\cup D_{i}[C^{1}]\cup D_{i}[C^{3}]\cup(\bigcup_{k\in[4]}D_{i}[C^{k},C^{k+1}]):i\in\mathcal{C}^{\prime\prime}\}$, say $M$. For all but at most one unused color in $\mathcal{C}^{\prime\prime}$, we can find a rainbow directed copy of $P_{3}$ with center in $Y$ and endpoints in $Z$, where $\{Y,Z\}=\{C^{1},C^{3}\}$. If some color $c\in\mathcal{C}^{\prime\prime}$ remains unused, then $D_{c}[C^{1},C^{3}]$ is nearly a complete bipartite graph and $D_{c}[C^{2}],D_{c}[C^{4}]$ are almost complete graphs. Then, by a similar argument to that in Theorem 6.5, $\mathcal{D}$ contains a transversal directed Hamilton cycle.
@@
We only consider the case that $|Y_{1}|\geq\gamma_{1}n$ and $|Y_{2}|<\gamma_{2}n$; other cases can be discussed similarly. Notice that vertices in $Y_{1}$ can be covered by a set of disjoint rainbow directed paths with centers in $Y_{1}$ and ending in $C^{4}\setminus(Y_{1}\cup Y_{2})$. In particular, those rainbow directed paths can be chosen such that none of them contains edges inside $\mathcal{D}[C^{4},C^{3}]$ or $\mathcal{D}[C^{1},C^{4}]$. Furthermore, vertices in $Y_{2}\setminus Y_{1}$ can be covered by a set of disjoint rainbow directed $P_{3}$ copies with centers in $Y_{2}\setminus Y_{1}$, starting at $C^{4}\setminus Y$, and using colors in $\mathcal{C}^{\prime}$.
In the digraph collection $\mathcal{D}[C^{4}\setminus Y]$, extend those rainbow directed paths or choose other disjoint rainbow directed paths into a set of disjoint maximal rainbow directed paths. Let $\mathbf{P}:=\{Q_{1},Q_{2},\ldots,Q_{t}\}$ be a set consisting of all disjoint rainbow paths in the above, each of which has at least one endpoint in $C^{4}\setminus Y$ and length $s_{i}\ (1\leq i\leq t)$ inside $C^{4}$. Let $\{e_{1},e_{2},\ldots,e_{r}\}$ be a set consisting of all disjoint rainbow edges in $\mathcal{D}[C^{4}\setminus Y,C^{3}]$, and let $\{\tilde{Q}_{1},\ldots,\tilde{Q}_{k}\}$ be a set of disjoint maximal rainbow directed $P_{3}$ in $\mathcal{D}^{\pm}[C^{4}\setminus Y,C^{1}]$ with centers in $C^{4}$ and endpoints in $C^{1}$, each of which is chosen while avoiding previously used vertices and colors.
@@
In this subcase, $\sum_{i=1}^{t}s_{i}+k+r<2\gamma n$. Assume that $V(Q_{i}[C^{4}]):=\{v_{s_{1}+\cdots+s_{i-1}+i},\ldots,v_{s_{1}+\cdots+s_{i}+i}\}$ for each $i\in[t]$. Let $w$ be an unused vertex in $C^{4}$ and $c_{1},c_{2}$ be two unused colors in $\mathcal{C}$. By the maximality of $\mathbf{P}$, in $D_{c_{1}}$ (resp. $D_{c_{2}}$), $w$ cannot be adjacent to the starting vertices of $Q_{i}[C^{4}]$ (resp. cannot be adjacent from the ending vertices of $Q_{i}[C^{4}]$) for all $i\in[t]$.